Circular No. 7181

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 K7 (LINEAR) An object discovered by LINEAR was reported as possibly cometary

by M. Elowitz, Lincoln Laboratory, with an apparent tail in p.a. ∼ 220^{o}. Confirmation of cometary activity has been reported by D. D. Balam (Victoria), who noted no tail but measured a 7′′ diffuse coma. The astrometry (with discovery observation provided below) is given on MPEC 1999-K37.

1999 UT  α_{2000}  δ_{2000}  m_{1}
May 24.34136  18^{h}27^{m}40.^{s}35  +19^{o}02′36′′.9  18.6

Preliminary parabolic orbital elements by B. G. Marsden:

T = 1999 Feb. 15.846 TT  ω = 100.^{o}487
Ω = 50.096 }2000.0
q = 2.27191 AU  i = 135.339

SUPERNOVAE 1999cg, 1999ch, 1999ci A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise

Observatory Optical Transients Search (cf. IAUC 6917, 7055) their discovery of three apparent supernovae on unfiltered CCD images taken with the Wise Observatory 1-m telescope:

SN  1999 UT  α_{2000}  δ_{2000}  R  Offset
1999cg Apr. 15 12^{h}43^{m}17.^{s}55 +76^{o}13′58′′.9 19.2  —
1999ch May 13 16 54 45.69 +39 59 13.9 19.8 0′′.5 west
1999ci May 15 14 52 12.70 +27 54 22.2 19.3  —

SNe 1999cg, 1999ch, and 1999ci are located in the fields of the galaxy clusters Abell 1607 (z = 0.136), Abell 2235 (z = 0.151), and Abell 1984 (z = 0.124), respectively. SNe 1999cg and 1999ci are superimposed on compact host galaxies, so no offsets can be determined from the images. SN 1999cg was confirmed on CCD images taken by Gal-Yam on May 12 with the same telescope, but it was invisible in images obtained in Mar. 20 (limiting mag R ∼ 22.0). SN 1999ch was confirmed on CCD images taken by Gal-Yam on May 16, but it was invisible in images obtained on 1998 May 2 (limiting mag R ∼ 21.0). SN 1999ci was confirmed on CCD images taken by Y. Lipkin on May 23 with the same telescope, but it was invisible in images obtained on 1998 May 27 (limiting mag R ∼ 21.0). Finding charts can be

obtained via http://wise-obs.tau.ac.il/~avishay/snlist.html.

1999 May 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green